# Phase error rate estimation in QKD with imperfect detectors

### v2

@@ VI.1 Recipe for computing key rates in the presence of basis-efficiency mismatch
Note that the above recipe is valid for for all detector models, and yields non-trivial results as long as one can suitably bound $\delta_{1},\delta_{2}$. In this work, we consider the canonical model of detectors in the next section, and then compute the values of $\delta_{1},\delta_{2}$ for this model Section VI.3.
VI.2 Detector Model
In this section, we specify the canonical model of Bob’s detectors (for active BB84) we use in this work. Let $\eta_{b_{i}},d_{b_{i}}$ denote the efficiency and dark count rate of Bob’s POVM corresponding to basis $b$, and bit $i$. We first define Bob’s double click POVM for basis $b\in\{Z,X\}$ to be $\Gamma^{(B)}_{(b,\text{dc})}=\sum_{N_{0},N_{1}=0}^{\infty}(1-(1-d_{b_{0}})(1-\eta_{b_{0}})^{N_{0}})(1-(1-d_{b_{1}})(1-\eta_{b_{1}})^{N_{1}})\outerproduct{N_{0},N_{1}}{N_{0},N_{1}}_{b}$ , where |$\outerproduct{N_{0},N_{1}}{N_{0},N_{1}}_{b}$ is the state with $N_{0}$ photons in the mode 1, and $N_{1}$ photons in mode 2, where the modes are defined with respect to basis $b$. For example, for polarisation-encoded BB84, |$\outerproduct{2,1}{2,1}_{Z}$ would signify the state with $2$ horizontally-polarised photons and $1$ vertically polarised photon. Recall that double clicks are mapped to single clicks randomly in our protocol. Thus, we can write Bob’s POVM elements as
@@ VI.2 Detector Model
$\Gamma_{(b.b),(\bot)}=$ $\mathrm{I}_{A}\otimes\sum_{N_{0},N_{1}=0}^{\infty}(1-d_{b_{0}})(1-d_{b_{1}})(1-\eta_{b_{0}})^{N_{0}}(1-\eta_{b_{1}})^{N_{1}}\outerproduct{N_{0},N_{1}}{N_{0},N_{1}}_{b}$ (58)
@@
VI.3 Computing bounds on $\delta_{1},\delta_{2}$
@@
VI.3.1 Active BB84 detection setup without any hardware modification
In this case the POVMs used by Alice and Bob are exactly given by Eq. 58. We construct the POVMs from Eqs. 55 and 54 in Section G.1. To bound $\delta_{1},\delta_{2}$, we use the fact that all POVMs are block-diagonal in the total photon number, and bound the $\infty$-norm of each block separately. Note that we can always treat the common value of loss in the detectors to be a part of the channel [52, Section III C]. This means that we pull out $(\max_{b,i}\{\eta_{b_{i}}\})$, and treat it as a part of the channel. (This is equivalent to giving the $\{\tilde{F},\mathrm{I}-\tilde{F}\}$ measurement to Eve.) This computation of $\delta_{1},\delta_{2}$ using the above steps is quite cumbersome, and is explained in Sections G.2 and G.3. Finally, we obtain
@@ VI.3.1 Active BB84 detection setup without any hardware modification
$d_{\text{max}}$ $=\max_{b\in\{X,Z\}}\{d_{b_{0}},d_{b_{1}}\}\leq d_{\text{det}}(1+\Delta_{\text{dc}}),\quad\text{ and }d_{\text{min}}=\min_{b\in\{X,Z\}}\{d_{b_{0}},d_{b_{1}}\}\geq d_{\text{det}}(1-\Delta_{\text{dc}}),$
$\eta_{\text{max}}$ $=\max_{b\in\{X,Z\}}\{\eta_{b_{0}},\eta_{b_{1}}\}\leq\eta_{\text{det}}(1+\Delta_{\eta}),\quad\text{ and }\eta_{\text{min}}=\min_{b\in\{X,Z\}}\{\eta_{b_{0}},\eta_{b_{1}}\}\geq\eta_{\text{det}}(1-\Delta_{\eta}).$
@@
VI.3.2 Random Swapping of 0 and 1 Detectors
@@ VI.4 Plots
VII Detector Side-Channels
The analysis presented so far assumes that the value of $\eta_{b_{i}},dc_{b_{i}}$ are fixed but unknown, and lie in some range given by Eq. 59. Thus, we have presented an analysis of Case 2 from Section I. This analysis allows us to drastically reduce the requirements on device characterisation: the proof technique is now robust to imperfect characterisation. However, physical implementations of QKD protocols are vulnerable to side-channel attacks where Eve can control, to a limited extent, the POVMs used. For example, by controlling the frequency, spatial mode [19] or arrival time [20] of the light, Eve can partially control the detector efficiencies and induce a suitable basis-efficiency mismatch. This is the scenario described by Case 3 from Section I.
We prove the security for such side-channels by expanding our detector model to account for all such spatio-temporal modes [13] as
$\Gamma=\bigoplus_{d}\Gamma_{\eta_{b_{i}}(d),d_{b_{i}}(d)},$ (66)
where $d$ denotes the spatio-temporal mode, and $\Gamma_{\eta_{b_{i}}(d),d_{b_{i}}(d)}$ denotes the POVM element corresponding to that mode. The detector has loss $\eta_{b_{i}}(d)$ for this mode, and a dark count rate of $d_{b_{i}}(d)$. The block-diagonal structure with respect to $d$ in the above equation reflects the fact that our model assumes no interference between any pair of spatio-temporal modes during the measurement process. Alternatively, one can equivalently consider a model where $d$ in Eq. 66 is interpreted as a classical flag, which Eve can use to choose detector efficiencies and dark count rates in their allowed ranges. While we only consider models of the above form in this work, our proof can accommodate a more complicated model with off-diagonal blocks as well (as long as one can suitably bound $\delta_{1},\delta_{2}$).
Due to the block-diagonal structure of the above POVM element Eq. 66, and the fact that $\delta_{1},\delta_{2}$ are $\infty$-norms which can be computed on each block-diagonal part separately, it is straightforward to see that our computation of $\delta_{1},\delta_{2}$ is directly applicable to the above scenario. In particular, the recipe from Section VI.1, and the computed key rates from Section VI.4 are valid for the scenario where Eve can control the value of $\eta_{b_{i}},d_{b_{i}}$ in the specified ranges (Eq. 59), via some extra spatio-temporal modes. Most importantly, our analysis does not depend on the number of such spatio-temporal modes. Thus, we are able to address scenarios where Eve has an arbitrary number of spatio-temporal modes, to induce (a bounded amount of) basis-efficiency mismatch in the detector.
This observation has great consequences, since the results of this work can now be used to prove security of QKD protocols in the finite-size regime in the presence of adversary-controlled basis-efficiency mismatch. Notably, this could not be accomplished with any proof technique before our work. Thus, we make a significant advancement towards the implementation security of QKD protocols.
VII.1 Independent and Correlated Detectors
Detectors are commonly modelled to act exactly the same way in each round. This assumption is made to simplify the security proof and is not physically motivated. For example, dead times and afterpulsing are non-IID effects that are present in realistic detectors. While we have presented our work for a model that assumes IID detector behaviour, we will now show that it can be easily extended to allow for non-IID behaviour of the detection setup as well; as long as the deviation is bounded.
@@ VII.1 Independent and Correlated Detectors
We first consider the case of independent detector behaviour in each round. In this case each round has a well-defined POVM that is independent of those in other rounds, i.e the POVM is tensor product with other rounds. Here we simply note that all the statistical claims used in our phase error estimation proofs (from Section C.2) remain true even if the POVM measurements are independent (but not IID). We comment on this and restate some of our lemmas for independent (but not identical) measurements in Section C.3. Moreover the Serfling statement (Lemma 2) does not assume any IID property of the input string. Thus, our bounds on the phase error rate remain unchanged as long as $\delta_{1},\delta_{2}$ can be bounded for the independent POVMs in each round. This deals with the case of independent detectors.
We believe a further extension to correlated detectors is possible using the ideas presented in Refs. [53, 8]. In this approach, a finite-correlation length $l_{c}$ of sources is assumed, and the security analysis is reduced to that of independent sources by implementing $l_{c}$ different QKD protocols that are “interleaved” with the rounds. That is, each successive round is considered to be a part of a one of the $l_{c}$ different QKD protocols. While we expect the above approach to be compatible with our work (by simply swapping sources with detectors), we do not formally do it in this work.
VIII Summary and Discussion

### v3

@@ VI.1 Recipe for computing key rates in the presence of basis-efficiency mismatch
VI.2 Assumptions
Note that the recipe is derived for the BB84 protocol (qubit or decoy-state) under the assumption that Alice’s source is perfect. That is, Alice sends either perfect qubit BB84 states, or perfectly phase-randomized weak coherent pulses for the decoy-state protocol. The recipe is valid for all active-choice detector models, and yields non-trivial results as long as one can suitably bound $\delta_{1},\delta_{2}$. For the explicit calculations in this work, we consider the canonical model of detectors in the next section, and then compute the values of $\delta_{1},\delta_{2}$ for this model in Section VI.4. Therefore, the bounds in Eq. 60 are derived assuming that Bob’s detector POVMs are given by Eq. 57 and characterized upto Eq. 59.
VI.3 Detector Model
In this section, we specify the canonical model of Bob’s detectors (for active BB84) we use in this work. Let $\eta_{b_{i}},d_{b_{i}}$ denote the efficiency and dark count rate of Bob’s POVM corresponding to basis $b$, and bit $i$. We first define Bob’s double click POVM for basis $b\in\{Z,X\}$ to be $\Gamma^{(B)}_{(b,\text{dc})}=\sum_{N_{0},N_{1}=0}^{\infty}(1-(1-d_{b_{0}})(1-\eta_{b_{0}})^{N_{0}})(1-(1-d_{b_{1}})(1-\eta_{b_{1}})^{N_{1}})\outerproduct{N_{0},N_{1}}{N_{0},N_{1}}_{b}$ , where |$\outerproduct{N_{0},N_{1}}{N_{0},N_{1}}_{b}$ is the state with $N_{0}$ photons in the mode 1, and $N_{1}$ photons in mode 2, where the modes are defined with respect to basis $b$. For example, for polarization-encoded BB84, |$\outerproduct{2,1}{2,1}_{Z}$ would signify the state with $2$ horizontally-polarised photons and $1$ vertically polarised photon. Recall that double clicks are mapped to single clicks randomly in our protocol. Thus, we can write Bob’s POVM elements as
@@ VI.3 Detector Model
$\Gamma_{(b,b),(\bot)}=$ $\mathrm{I}_{A}\otimes\sum_{N_{0},N_{1}=0}^{\infty}(1-d_{b_{0}})(1-d_{b_{1}})(1-\eta_{b_{0}})^{N_{0}}(1-\eta_{b_{1}})^{N_{1}}\outerproduct{N_{0},N_{1}}{N_{0},N_{1}}_{b}$ (58)
@@
VI.4 Computing bounds on $\delta_{1},\delta_{2}$
@@
VI.4.1 Active BB84 detection setup without any hardware modification
In this case the POVMs used by Alice and Bob are exactly given by Eq. 58. We construct the POVMs from Eqs. 55 and 54 in Section G.1. To bound $\delta_{1},\delta_{2}$, we use the fact that all POVMs are block-diagonal in the total photon number, and bound the $\infty$-norm of each block separately. Note that we can always treat the common value of loss in the detectors to be a part of the channel [59, Section III C]. This means that we pull out $(\max_{b,i}\{\eta_{b_{i}}\})$, and treat it as a part of the channel. (This is equivalent to giving the $\{\tilde{F},\mathrm{I}-\tilde{F}\}$ measurement to Eve.) This computation of $\delta_{1},\delta_{2}$ using the above steps is quite cumbersome, and is explained in Sections G.2 and G.3. Finally, we obtain
@@ VI.4.1 Active BB84 detection setup without any hardware modification
$d_{\text{max}}$ $=\max\{d_{X_{0}},d_{X_{1}},d_{Z_{0}},d_{Z_{1}}\}\leq d_{\text{det}}(1+\Delta_{\text{dc}}),\quad\text{ and }d_{\text{min}}=\min\{d_{X_{0}},d_{X_{1}},d_{Z_{0}},d_{Z_{1}}\}\geq d_{\text{det}}(1-\Delta_{\text{dc}}),$
$\eta_{\text{max}}$ $=\max\{\eta_{X_{0}},\eta_{X_{1}},\eta_{Z_{0}},\eta_{Z_{1}}\}\leq\eta_{\text{det}}(1+\Delta_{\eta}),\quad\text{ and }\eta_{\text{min}}=\min\{\eta_{X_{0}},\eta_{X_{1}},\eta_{Z_{0}},\eta_{Z_{1}}\}\geq\eta_{\text{det}}(1-\Delta_{\eta}).$
@@
VI.4.2 Random Swapping of 0 and 1 Detectors
@@ VI.5 Plots
  4. 4.
    In Fig. 7, we plot the finite size key rate against both detector characterization parameters $\Delta_{\eta},\Delta_{\text{dc}}$ independently. We find that both the parameters $\Delta_{\eta},\Delta_{\text{dc}}$ lead to comparable penalties in the key rate, although $\Delta_{\text{dc}}$ penalizes the key rate less than $\Delta_{\eta}$. Note that the values for $\delta_{1},\delta_{2}$ also depend on the dark count rate $d_{\text{det}}$.
@@
Figure 7: Finite-size key rates for various values of $\Delta_{\eta},\Delta_{\text{dc}}$ for the decoy-state BB84 protocol, for $N_{\text{tot}}=10^{12}$ and $25$dB loss. We find that both $\Delta_{\eta},\Delta_{\text{dc}}$ have comparable impact on the keyrate, although $\Delta_{\text{dc}}$ penalizes the key rate less than $\Delta_{\eta}$. The above plot is interpolated from key rate calculations of $2500$ points, and the detectors are not swapped randomly.
@@
We end this this section by considering a scenario where the detector behavior is independent (but not identical) in each round. In this case each round has a well-defined POVM that is independent of those in other rounds, i.e the POVM is tensor product with other rounds. Here we simply note that all the statistical claims used in our phase error estimation proofs (from Section C.2) remain true even if the POVM measurements are independent (but not IID). We comment on this and restate some of our lemmas for independent (but not identical) measurements in Section C.3. Moreover the Serfling statement (Lemma 2) does not assume any IID property of the input string. Thus, our bounds on the phase error rate remain unchanged as long as $\delta_{1},\delta_{2}$ can be bounded for the independent POVMs in each round. Such a scenario is of practical importance, as detection setups are never perfectly IID [60, Fig. 3(a)]. Another important practical consideration is that of correlated detectors, which we now discuss in the next section.
VII Application to Correlated effects
We now turn our attention to detectors exhibiting correlated behavior across the rounds. We begin by formalizing our model for such detectors. Let the outcome of round $i$ for Bob be denoted by $k_{i}$, and define $k_{i}^{j}$ as the sequence of outcomes from round $i$ to round $j$ ($i\leq j$). Recall that $\bot$ corresponds to the no-detect outcome. Correlated effects such as afterpulsing and detector dead times can be modeled by allowing the POVM used in later rounds to depend on the outcome — in particular, detection events — of previous rounds. This is the scenario we are interested in. More formally, in the $i$th round, the POVM used for the measurement is given by $\{\Gamma_{k_{i}}^{(k^{i-1}_{i-l_{c}})}\}$, where $l_{c}$ denotes the correlation length. Note that $l_{c}$ here denotes the correlation length in units of the time slots (i.e, the unit is the time between successive measurements in the QKD protocol). We use the convention that when all $k_{i-l_{c}}^{i-1}=\bot s$, then the superscript can be omitted.
Our approach involves a protocol modification in which Alice and Bob retain only those rounds where previous $l_{c}$ rounds involved no-click events. All other rounds are ‘rejected’, i.e thrown away. In our terminology, a round can be rejected in this manner (due to the specific postprocessing described above), but may also be discarded later based on its detection outcome -— for instance, if it results in a no-detection event. The idea is that after this postprocessing, the remaining rounds can be thought of as all being measured using uncorrelated POVM $\{\Gamma_{k_{i}}\}$.
The key idea is to introduce a generalized filtering step (analogous to $\tilde{F}$ in the round-by-round case) that performs the minimum amount of measurements necessary, and uses the outcomes to reject rounds based on the above postprocessing. Once the rounds to be rejected are fixed, the remaining rounds can be measured fully using uncorrelated POVMs, and the analysis on these rounds can follow the same approach as the usual EUR analysis.
However, note that the postprocessing described above depends on the detection events, which in turn depend on the choice of basis, and thus can only be determined after basis choice. Thus, in general, a basis choice needs to be made before rounds can be rejected. This is the core difficulty in proving security under detector correlations within the EUR framework, as the EUR statement must be applied to the state before choosing the basis.
We can now present a two-step proof sketch to address the above issue (we elaborate on each step in later subsections):
  1. 1.
    Assuming the detectors have no loss or dark counts, the rejection step can be performed in a basis-independent manner (on the rounds which will not be rejected). Thus, the EUR statement can be applied to those rounds by completing the measurements later. We explicitly construct this procedure in Section VII.1, and show that is rejects the correct rounds (see Lemma 5).
  2. 2.
    We lift the assumption (of no loss and no dark counts) made in the first step above using techniques from Ref. [61]. Intuitively, Ref. [61] allows one to incorporate the effects of basis-dependent dark counts and efficiencies at some cost (to be discussed in Section VII.2). Thus, this step can be done first, reducing the problem to the analysis described in the earlier step.
VII.1 Perfect correlated detectors
The intuition behind the first step relies on the following two observations. First, detections in the rounds that will not be rejected can be inferred from the photon number of the incoming state, independent of basis choice. This is because the previous rounds have had no clicks, so the POVM used in the current is the uncorrelated, perfect one. Moreover, due to the block-diagonal nature of the POVMs, this photon number measurement does not affect the measurement statistics. Second, the EUR statement is not used on the rounds that are rejected. Thus, we can are allowed to complete the measurement on these rounds.
Lemma 5.
Consider the state $\rho_{Q^{n}}$, and let the $i$th subsystem be measured using POVM $\{\Gamma_{k_{i}}^{(k^{i-1}_{i-l_{c}})}\}$, where $k_{i}$ denotes the outcome of the $i$th measurement, and $k^{i-1}_{i-l_{c}}$ denotes the string of outcomes in the previous $l_{c}$ rounds. Suppose that $\{\Gamma_{k_{i}}\}$ (corresponding to $k^{i-1}_{i-l_{c}}=\bot s)$ is such that it can be described by a two-step measurement, where the first measurement $\{\tilde{F},\mathrm{I}-\tilde{F}\}$ determines the detect vs no-detect, followed by a second step measurement. Consider the state obtained after rejecting all rounds $i$ for which a detection occurred in the previous $l_{c}$ rounds. Then this state can be obtained via a procedure that only performs the $\{\tilde{F},\mathrm{I}-\tilde{F}\}$ measurements on the rounds that are not rejected. This can be formalized in the following lemma.
Proof.
For simplicity, we assume a correlation length $l_{c}$ of $1$. The extension to larger, finite correlation lengths is straightforward. We will prove the required claim by explicit construction of the procedure. We will prove it sequentially, from round $1$ to round $n$.
First note that round $1$ will never be rejected. The $\{\tilde{F},\mathrm{I}-\tilde{F}\}$ measurement in round $1$ directly tells us whether or not the round will result in a detection event. Thus, this tells us whether or not the next round ($i=2$) will be rejected. (Note that this measurement was performed without choosing a measurement basis for round $1$.)
For round $i\geq 3$, assume rounds $1$ to $i{-}1$ already have definite reject/no-reject tags, with only the $\{\tilde{F},\mathrm{I}-\tilde{F}\}$ measurement being performed on the rounds that are not rejected. Then, the round $i$ can be given the correct reject/no-reject tag as follows.
  1. 1.
    To determine whether round $i$ should be kept, check the reject/no-reject tag of round $i-1$ (which is guaranteed to be correct).
  2. 2.
    If round $i-1$ is to be rejected, complete the measurement on it. This determines whether or not there is a detection event in this round, which inturn determines whether or not round $i$ will be rejected.
  3. 3.
    If round $i-1$ is not rejected, then round $i-2$ must have had no detection (since otherwise the round $i-1$ would be rejected). This means that the POVM to be used in round $i-1$ is the uncorrelated one ($\{\Gamma_{k_{i-1}}\}$). Thus, measuring $\{\tilde{F},\mathrm{I}-\tilde{F}\}$ directly tells us whether or not round $i-1$ will result in a detection event. This, in turn, determines whether or not round $i$ will be rejected.
By construction, the above procedure only implements the $\{\tilde{F},\mathrm{I}-\tilde{F}\}$ on the rounds that are not rejected, and results in the correct rounds being rejected. This concludes our proof.
KeepDiscardKeepRound $i{-}2$Round $i{-}1$Round $i$DetectNo-detectNo-detectEarlierroundsLaterrounds Photon-number measurement Complete measurement Photon-number measurement
Figure 8: Illustration of the process to determine the rounds that will be rejected based on prior detection events. The rounds that are not rejected are only measured using $\{\tilde{F},\mathrm{I}-\tilde{F}\}$, which determines the detect vs no-detect outcome. They will have the rest of the measurement, including basis choice, completed at a later stage. The rounds that are rejected have their entire measurements completed in order to determine the detection event.
VII.2 Reduction to perfect correlated detectors
As mentioned earlier, the applicability of Lemma 5 crucially relies on the fact that the detectors are perfect threshold detectors – except for correlations that one detect event might trigger other detect events within the correlation length. However, as argued in the rest of this paper, this is not a practical assumption. To resolve this issue, we use the results of Ref. [61].
Ref. [61] constructs a basis-choice independent reduction from a detection setup with detectors with different dark counts and loss, to a detection setup with perfect detectors. It shows that the imperfect detection setup can be treated as a basis-choice independent channel (similar to the basis-independent filter $\tilde{F}$) followed by the perfect detection setup. To do this, it crucially relies on the flag-state squasher [13]. The intuition is that the flag-state squasher makes it possible to ‘give’ Eve full information about multi-photon signals by introducing a classical flag space. Crucially, this flag space can also then be used to transfer imperfections such as loss and dark counts to Eve. This is argued formally in [61, Theorem 1 and 2].
The main open task in our approach outlined in Section VII is the use of the flag-state squasher in an EUR based security proof. In the typical usage of the flag-state squasher, one considers a single round state, and the weight of the state in the flag space is bounded through suitable methods to prevent trivial results (since the entire flagged state is revealed to Eve). In the EUR context, there no single round state. Instead, the natural alternative is to instead bound the number of rounds in the flag space. We believe that the usual methods of bounding the weight of the state along with suitable concentration inequalities will suffice for this task. However, we emphasize that in this work, we provide only a sketch of the proof for handling correlated detectors, leaving a detailed analysis to future work — a non-trivial task. In the next section, we shift our focus to detector side-channels, and in particular, we will see how our methods naturally address certain detector side-channels as a by-product.
VIII Detector Side-Channels
The analysis presented so far assumes that the detector behavior, while possibly varying between rounds or exhibiting correlations, is described by a single mode characterized by bounded loss and dark count rates $\eta_{b_{i}}$ and $d_{b_{i}}$, as specified in Eq. 59. Thus, we have presented an analysis of Case 2 from Section I. This analysis allows us to drastically reduce the requirements on device characterization: the proof technique is now robust to imperfect characterization. However, physical implementations of QKD protocols are vulnerable to side-channel attacks where Eve can control, to a limited extent, the POVMs used. For example, by controlling the frequency, spatial mode [62, 19] or arrival time [20] of the light, Eve can partially choose the detector efficiencies and induce a suitable basis-efficiency mismatch. This is the scenario described by Case 3 from Section I.
While our proof technique advances the theory to the point where this case can be handled in principle, a complete analysis first requires the physical modeling of multi-mode detectors, which remains an open problem. In this section, we outline how the results of this work can be applied to a simple multi-mode model.
We expand our detector model (and Bob’s Hilbert space) to account for spatio-temporal modes [13] as
$\Gamma^{\text{multi}}_{(b_{A},b_{B}),(k)}=\bigoplus_{\mathbf{d}}\Gamma_{(b_{A},b_{B}),(k)}(\{{\eta_{b_{i}}(\mathbf{d}),d_{b_{i}}(\mathbf{d})}\}),$ (66)
where $\mathbf{d}$ denotes the spatio-temporal mode, and $\Gamma_{(b_{A},b_{B}),(k)}(\{{\eta_{b_{i}}(\mathbf{d}),d_{b_{i}}(\mathbf{d})}\})$ denotes the single-mode POVM element corresponding to that mode, and is given by Eq. 58. The multi-mode detector has loss $\eta_{b_{i}}(\mathbf{d})$ for this mode, and a dark count rate of $d_{b_{i}}(\mathbf{d})$.
The block-diagonal structure with respect to $\mathbf{d}$ in the above equation reflects the fact that our model assumes no interference between any pair of spatio-temporal modes during the measurement process. In particular, it captures the possibility that an adversary may exploit different times-of-arrival, frequencies, or angles of incidence to attack the system, provided that each instance corresponds to a definite spatio-temporal mode and no coherent superpositions across modes are used. Even with this limitation, the model protects against a wide range of known classical side-channel attacks. For instance, the time-shift attack [17, 20] is fully captured within this model, as it simply corresponds to Eve selecting different times-of-arrival to exploit the time-dependent efficiency mismatch of the gated detectors. Thus, the block-diagonal model represents a first step toward a more complete analysis of realistic side-channels. This perspective also captures other potential attack strategies, such as modifying the temperature of the detection setup.
Remark 15.
We stress that our results in this subsection should be interpreted within the context of this model, and may not accurately describe the physical reality of multi-mode detectors. Nevertheless, while we only consider models of the above form in this work, our proof provides a framework to accommodate more complicated models of multi-mode detectors with off-diagonal blocks, as long as one can suitably bound $\delta_{1},\delta_{2}$. In general, this would require a model of the detectors, and characterization of the detectors over all the modes. For examples of such attempts to experimentally characterize all the modes, see Ref. [62, 19].
Due to the block-diagonal structure of the above POVM element Eq. 66, and the fact that $\delta_{1},\delta_{2}$ are $\infty$-norms which can be computed on each block-diagonal part separately, it is straightforward to see that our computation of $\delta_{1},\delta_{2}$ is directly applicable to the above scenario. To see this, note that our metrics are obtained by first constructing POVMs corresponding to a multi-step measurement process, as outlined in Section VI.1. This construction preserves the block-diagonal structure of Eq. 66. Thus, if $\delta_{1}(\{{\eta_{b_{i}}(\mathbf{d}),d_{b_{i}}(\mathbf{d})}\}),\delta_{2}(\{{\eta_{b_{i}}(\mathbf{d}),d_{b_{i}}(\mathbf{d})}\})$ are the values of these metrics computed according to Eq. 56, for the appropriate single-mode POVMs, then the metrics for the multi-mode case are given by
$\delta_{1}^{\text{multi}}=$ $\max_{\mathbf{d}}\delta_{1}(\{{\eta_{b_{i}}(\mathbf{d}),d_{b_{i}}(\mathbf{d})}\}),$ (67)
$\delta_{2}^{\text{multi}}=$ $\max_{\mathbf{d}}\delta_{2}(\{{\eta_{b_{i}}(\mathbf{d}),d_{b_{i}}(\mathbf{d})}\}).$
If the values of $\{{\eta_{b_{i}}(\mathbf{d}),d_{b_{i}}(\mathbf{d})}\}$ are characterized and satisfy Eq. 59 for all $\mathbf{d}$, then Eq. 67 is exactly the same as the computation as in Step (5) of Section VI.1 (which corresponds to computing $\delta_{1},\delta_{2}$ for Case 2 from Section I).
This means that the recipe from Section VI.1, and the computed key rates from Section VI.5 are valid for the scenario where Eve can choose the value of $\eta_{b_{i}},d_{b_{i}}$ in the specified ranges (Eq. 59), via some extra spatio-temporal modes. Most importantly, our analysis does not depend on the number of such spatio-temporal modes. Thus, we are able to address scenarios where Eve has an arbitrary number of spatio-temporal modes, to induce (a bounded amount of) basis-efficiency mismatch in the detector.
Remark 16.
As discussed above, our methods are such that allowing Eve to choose the detector parameters within the characterized range yields the same key rate as having fixed detector parameters that are characterized within the same range. However, this observation need not be fundamental, and may be a consequence of the proof technique used in this work. This is because intuitively, we expect scenarios where Eve cannot choose the detectors to lead to higher key rates that scenarios where she can. Nevertheless, while we do not know of a physical mechanism by which Eve can choose dark count rates, we allow Eve to choose them along with the detection efficiency.
We have picked this model for its theoretical simplicity. However, more realistic models such as the one introduced in [12, Section 3] can also be analysed with the results in this work. In that case, the computation of $\delta_{1}$ and $\delta_{2}$ would constitute a more involved version of our current computations described in Appendix G. Specifically, Eq. 121 would need to be modified with a different choice of operator $P$. We note that the basis dependent filters for this model are still block-diagonal in the total number of photons $n$ across all modes. Moreover, as $n$ increases, the filtering operators approach the identity operator (since the probability of detect approaches $1$). Thus, we expect $\delta_{1}$ and $\delta_{2}$ to depend on the $n\leq 1$ blocks. If this monotonicity can be rigorously proven, then the $n\leq 1$ block contributions to $\delta_{1}$ and $\delta_{2}$ can even be computed numerically.
Finally we note that this work does not apply to all detector side-channels. For instance, our model does not fit Trojan horse attacks [63]. Moreover, some blinding attacks on detectors [64] lead to complete knowledge of Bob’s detection events to Eve. In this case, our methods naturally lead to trivial key rates, since no key generation is possible.
IX Summary and Discussion